MARR is 5%

what is the incremental rate and what alternative should be chosen.

YEAR A B
0 -10000 -15000
1-5 3200 4500
Answer #1

incremental initial cost (B-A) = -15000 - (-10000) = -5000

incremental annual benefit (B-A) = 4500 - 3200 = 1300

Let i%, be incremental IRR, then

1300 * (P/A,i%,5) = 5000

(P/A,i%,5) = 5000 / 1300 = 3.846153

using trail and error method

When I = 5%, value of (P/A,i%,5) = 4.329477

When I = 9%, value of (P/A,i%,5) = 3.889651

When I = 10%, value of (P/A,i%,5) = 3.790787

using interpolation

I = 9% + (3.889651 -0 3.846153) / (3.889651 - 3.790787)*(10%-9%)

I = 9% + 0.4399%

I = 9.4% (Approx)

As incremental IRR > MARR, we should select option B